OpenStudy (anonymous):

solve the equation -7(2x-1)=35 x=???

OpenStudy (lgbasallote):

divide both sides by -7 first what do you get

OpenStudy (anonymous):

-5? @lgbasallote

OpenStudy (anonymous):

On the left hand side what you get???

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

There is a term too with it..

OpenStudy (anonymous):

? dont know

OpenStudy (lgbasallote):

what is \[-7(2x - 1) \div -7?\]

jimthompson5910 (jim_thompson5910):

That -5 is what is left on the right side after division

jimthompson5910 (jim_thompson5910):

So -7(2x-1)=35 becomes 1(2x-1) = -5 which turns into 2x-1 = -5

jimthompson5910 (jim_thompson5910):

Remember that -7/(-7) is 1 (since dividing any number but zero by itself gives you 1)

OpenStudy (anonymous):

1.8? haha sorry I'm so confused

jimthompson5910 (jim_thompson5910):

Do you see how I got 2x-1 = -5 ?

OpenStudy (anonymous):

yea kinda haha

jimthompson5910 (jim_thompson5910):

If not, then go back over it and ask about any piece that's getting you stuck If so, then your task is to solve 2x-1 = -5 So your next step is what?

OpenStudy (anonymous):

-2 ?

jimthompson5910 (jim_thompson5910):

You got it

jimthompson5910 (jim_thompson5910):

Nice work

OpenStudy (anonymous):

Thank you!!!

jimthompson5910 (jim_thompson5910):

you're welcome